# Bayes optimal learning of attention-indexed models

### v2

@@ C.3 Output Channel Computation
$I_{\rm out}(q)=\left[\int dy\int\prod_{u,\ell,a\leq b}dh^{(\ell)\;u}_{ab}\,P\left(\{h^{(\ell)\;u}_{ab}\}_{ab}\right)\prod_{u,\ell}P_{\rm out}\left(y|\Bigr\{\frac{h^{(\ell)\;u}_{ab}}{\sqrt{2-\delta_{ab}}}\Bigr\}_{ab}\right)\right]^{n}.$ (97)
@@
$\mathbb{E}[h^{(\ell)}_{ab}]=0\qquad\operatorname{Cov}_{x^{\mu}}\left(h^{(\ell)\;u}_{a\leq b},h^{(k)\;v}_{c\leq d}\right)=\frac{1}{d}\left[2\;\delta_{ac}\delta_{bd}\right]\operatorname{Tr}\left(S^{u}_{\ell}S^{v}_{k}\right)=\left[2\;\delta_{ac}\delta_{bd}\right]Q^{uv}_{\ell k}$ (98)
@@
$\frac{1}{2}\sum_{u,v=0}^{m}\sum_{a\leq b,c\leq d}\sum_{\ell,k}\biggl(h^{(\ell)\;u}_{ab}\biggr)\left(\Sigma_{h}^{-1}\right)^{uv,\ell k}_{ab,cd}\biggl(h^{(k)\;v}_{cd}\biggr)$ (99)
@@
$I_{\rm out}(q)=\left[\int dy\int\prod_{u,\ell,a\leq b}dh^{(\ell)\;u}_{ab}\,\exp\left(-\frac{1}{2}\sum_{u,v=0}^{m}\sum_{a,b,c,d}\sum_{\ell,k}h^{(\ell)\;u}_{ab}\left(\Sigma_{h}^{-1}\right)^{uv,pq}_{a\leq b,c\leq d}h^{(k)\;v}_{cd}\right)\prod_{u,\ell}P_{\rm out}\left(y|\Bigr\{\frac{h^{(\ell)\;u}_{ab}}{\sqrt{2-\delta_{ab}}}\Bigr\}_{ab}\right)\right]^{n}$
@@
$\left(\Sigma_{h}\right)^{uv,\ell k}_{a\leq b,c\leq d}\;=\;\left[2\;\delta_{ac}\delta_{bd}\right]\,Q^{uv}_{\ell k}\;=\;\left[2\;\delta_{ac}\delta_{bd}\right]\,\Bigl[(Q_{\ell k}-q_{\ell k})\,\delta_{uv}+q_{\ell k}\Bigr]\,$, (101)
The inverse of the covariance matrix is given by:
$\bigl[(\Sigma^{-1})_{ab}\bigr]^{\,u\ell,\,vq}=\frac{1}{2}\;\Bigl[\bigl[(Q-q)^{-1}\bigr]_{\ell k}\,\delta_{uv}\;-\;\sum_{s=1}^{L}\sum_{t=1}^{L}\bigl[(Q-q)^{-1}\bigr]_{\ell s}\,q_{st}\,\bigl[Q^{-1}\bigr]_{tk}\Bigr]$ (102)
Introducing $M_{\ell}:=\sum_{u=0}^{m}h^{(\ell)\;u}_{ab}$, we can rewrite the Gaussian exponent in Eq. (100) as:
$\frac{1}{4}\biggl[\sum_{u,\ell,k}h^{(\ell)\;u}_{ab}\,\bigl[(Q-q)^{-1}\bigr]_{\ell k}\,h^{(k)\;u}_{ab}-\sum_{\ell,k}M_{\ell}\;\underbrace{\Bigl(\sum_{s,t}\bigl[(Q-q)^{-1}\bigr]_{\ell s}\,q_{st}\,\bigl[Q^{-1}\bigr]_{tk}\Bigr)}_{=:C_{\ell k}}\;M_{k}\biggr].$ (103)
We now introduce L auxiliary Gaussian variables for this token pair, $\eta^{(\ell)}_{ab}\sim\mathcal{N}(0,1)$ ($\ell=1,\dots,L$), via
$\exp\!\Bigl(\tfrac{1}{4}\sum_{\ell,k}M^{(\ell)}_{ab}C_{\ell k}M^{(k)}_{ab}\Bigr)=\int\!\prod_{s=1}^{L}\frac{d\eta^{(s)}_{ab}}{\sqrt{2\pi}}\,\exp\!\Bigl(-\tfrac{1}{2}\!\sum_{s}(\eta^{(s)}_{ab})^{2}+\tfrac{1}{2}\!\sum_{\ell,k}\eta^{(\ell)}_{ab}\,\sqrt{C}_{\ell k}\,M^{(k)}_{ab}\Bigr)$, (104)
where the square root is intended over the spectrum of the matrix. After some algebra, we can manipulate the Gaussian exponent in the following way:
$\frac{1}{4}\sum_{\ell,k}\Bigl(h^{(\ell)\;u}_{ab}-\sum_{s=1}^{L}\sqrt{2q}_{\ell s}\;\eta^{\,(s)}_{ab}\Bigr)\bigl[(Q-q)^{-1}\bigr]_{\ell k}\Bigl(h^{(k)\;u}_{ab}-\sum_{t=1}^{L}\sqrt{2q}_{kt}\;\eta^{\,(t)}_{ab}\Bigr)-\frac{1}{2}\sum_{r}(\eta^{(r)}_{ab})^{2}.$ (105)
We finally recognize:
$\frac{1}{2}\sum_{\ell=1}^{L}\sum_{k=1}^{L}\bigl(h^{(\ell)\;u}_{ab}-\omega^{(\ell)}_{ab}\bigr)\;[V^{-1}_{ab}]_{\ell k}\;\bigl(h^{(k)\;u}_{ab}-\omega^{(k)}_{ab}\bigr)\;-\frac{1}{2}\sum_{r}(\eta^{(r)}_{ab})^{2}$ (106)
with
$\omega_{ab}^{(\ell)}=\sum_{k=1}^{L}\sqrt{2q}_{\ell k}\,\eta_{ab}^{(k)}\,,\qquad V^{(\ell k)}=2(Q_{\ell k}-q_{\ell k})$ (107)
Compatibly with the results of [Aubin2018, troiani2024fundamental]. Hence we finally recognize a Gaussian term of the form:
$\int\prod_{a\leq b}^{T}d^{L}h_{ab}\,\mathcal{N}\left(h_{ab};\omega_{ab};V_{ab}\right)P_{\rm out}\left(y\mid h^{u}_{ab}\right)$ (108)
Since replicas are decoupled through $\eta$, we can finally write out output channel term as:
@@
$I_{\rm out}(q)=\left[\int dy\int\prod_{a\leq b,\ell}\frac{d\eta_{ab}^{(\ell)}}{\sqrt{2\pi}}e^{-\frac{(\eta_{ab}^{(\ell)})^{2}}{2}}\left(\int\prod_{a\leq b}^{T}d^{L}h_{ab}\mathcal{N}\left(h_{ab};\omega_{ab},V_{ab}\right)P_{\rm out}\left(y\mid\{h_{ab}^{(\ell)}\}_{\ell=1}^{L}\right)\right)\right]^{n(m+1)}$
Which, as for the prior term, we can write as:
$I_{\rm out}(q)=\left[\int dy\int\mathcal{D}\eta\left(\mathcal{Z}_{\rm out}(y,\omega,V)\right)\right]^{n(m+1)}$ (110)
and:
$\mathcal{Z}_{\rm out}(y,\omega,V)=\int\prod_{a\leq b}d^{L}h_{ab}\mathcal{N}\left(h_{ab};\omega_{ab},V_{ab}\right)P_{\text{out}}\left(y\mid\{h_{ab}^{(\ell)}\}_{\ell=1}^{L}\right)$ (111)
Where we defined the measure over the auxiliary variables as
$\mathcal{D}\eta=\prod_{\ell=1}^{L}\mathcal{D}\eta^{(\ell)}\;=\;\prod_{\ell=1}^{L}\prod_{a\leq b}\frac{d\eta_{ab}^{(\ell)}}{\sqrt{2\pi}}\,\exp\!\Bigl[-\frac{(\eta_{ab}^{(\ell)})^{2}}{2}\Bigr]\,.$ (112)
Expanding Eq. (110) for small number of replicas $m$ we get:
@@
Recalling the state equations found in Eq. (100) we get:
$\hat{q}_{\ell k}=4\alpha\int D\eta\int dy\,\left(\frac{\partial\mathcal{Z}_{\rm out}(y,\omega,V)}{\partial q_{\ell k}}\left(1+\ln\mathcal{Z}_{\rm out}(y,\omega,V)\right)\right)\,.$ (114)
Now, since we know that
$\frac{\partial\omega_{ab}^{(r)}}{\partial\eta_{ab}^{(k)}}=(\sqrt{2q})_{rk}$ (115)
for each token pair $(a,b)$, we define the denoising function:
$(g_{\rm out}(y,\omega,V))_{ab}^{(\ell)}=\frac{\partial}{\partial\omega_{ab}^{(\ell)}}\ln\mathcal{Z}_{\rm out}(y,\omega,V)=\frac{\partial\mathcal{Z}_{\rm out}(y,\omega,V)}{\partial\omega_{ab}^{(\ell)}}\mathcal{Z}_{\rm out}(y,\omega,V)^{-1}\,.$ (116)
So, using the chain rule and the definition (116):
$\partial_{\eta_{ab}^{(k)}}\mathcal{Z}_{\mathrm{out}}=\sum_{r=1}^{L}\frac{\partial\mathcal{Z}_{\mathrm{out}}}{\partial\omega_{ab}^{(r)}}\frac{\partial\omega_{ab}^{(r)}}{\partial\eta_{ab}^{(k)}}=Z_{\mathrm{out}}(y,\omega,V)\sum_{r=1}^{L}g_{ab}^{(r)}(y,\omega,V)\,\mathcal{\;}(\sqrt{2q})_{rk}.$
We can write the derivative explicitly:
$\frac{\partial\mathcal{Z}_{\mathrm{out}}}{\partial q_{\ell k}}=\sum_{a\leq b}\sum_{r=1}^{L}\frac{\partial\mathcal{Z}_{\mathrm{out}}}{\partial\omega_{ab}^{(r)}}\frac{\partial\omega_{ab}^{(r)}}{\partial q_{\ell k}}\;+\;\frac{\partial\mathcal{Z}_{\mathrm{out}}}{\partial V_{\ell k}}\frac{\partial V_{\ell k}}{\partial q_{\ell k}}.$ (117)
Regarding the first term, one can show that:
$\frac{\partial\omega_{ab}^{(r)}}{\partial q_{\ell k}}=(\sqrt{2q})^{\,-1}_{\ell k}\,\omega_{ab}^{(r)}\;+\;\text{(sym.\ in $\ell,k$)}.$ (118)
Multiplying by $\partial_{\omega_{ab}^{(r)}}\mathcal{Z}_{\mathrm{out}}=g_{ab}^{(r)}\,\mathcal{Z}_{\mathrm{out}}$ and using $\omega_{ab}^{(r)}=\sum_{s}(\sqrt{2q})_{rs}\eta_{ab}^{(s)}$ gives
$\frac{1}{2}\sum_{a\leq b}\bigl[\eta_{ab}^{(k)}\,g_{ab}^{(\ell)}+\eta_{ab}^{(\ell)}\,g_{ab}^{(k)}\bigr]\,\mathcal{Z}_{\mathrm{out}}\!.$ (119)
Regarding the second piece contribution, we first notice $V_{\ell k}=2(Q_{\ell k}-q_{\ell k})$, giving us $\partial_{q_{\ell k}}V_{\ell k}=-2.$ A straightforward but lengthy calculation shows that the term coming from $\partial\mathcal{Z}_{\mathrm{out}}/\partial V_{\ell k}$ exactly cancels the $\nicefrac{{1}}{{2}}$ factor above. Finally, combining the two pieces yields
$\frac{\partial\mathcal{Z}_{\mathrm{out}}}{\partial q_{\ell k}}=\sum_{a\leq b}\Bigl[\eta_{ab}^{(k)}\,g_{ab}^{(\ell)}+\eta_{ab}^{(\ell)}\,g_{ab}^{(k)}\Bigr]\,\mathcal{Z}_{\mathrm{out}}$ (120)
which, by the definition of the denoising function and the means $\omega_{ab}^{(\ell)}$ can be manipulated into:
$\frac{\partial\mathcal{Z}_{\mathrm{out}}}{\partial q_{\ell k}}=\frac{1}{2\,q_{\ell k}}\sum_{a\leq b}\Bigl[\eta_{ab}^{(k)}\,\partial_{\eta_{ab}^{(\ell)}}\mathcal{Z}_{\mathrm{out}}+\eta_{ab}^{(\ell)}\,\partial_{\eta_{ab}^{(k)}}\mathcal{Z}_{\mathrm{out}}\Bigr].$ (121)
Now, recalling that:
$\widehat{q}_{\ell k}=4\alpha\int\!\mathcal{D}\eta\,dy\;\bigl[1+\ln\mathcal{Z}_{\mathrm{out}}\bigr]\;\partial_{q_{\ell k}}\mathcal{Z}_{\mathrm{out}}.$ (122)
Inserting (121) leads:
@@
$\widehat{q}_{\ell k}$ $=\frac{2\alpha}{q_{\ell k}}\sum_{a\leq b}\!\int\!\mathcal{D}\eta\,dy\;\bigl[1+\ln\mathcal{Z}_{\mathrm{out}}\bigr]\bigl[\eta_{ab}^{(k)}\,\partial_{\eta_{ab}^{(\ell)}}\mathcal{Z}_{\mathrm{out}}+\eta_{ab}^{(\ell)}\,\partial_{\eta_{ab}^{(k)}}\mathcal{Z}_{\mathrm{out}}\bigr].$ (123)
Using $\int\!\mathcal{D}\eta\;\eta F=\int\!\mathcal{D}\eta\;\partial_{\eta}F$ on both terms, i.e.
$\int\!\mathcal{D}\eta\,\eta_{ab}^{(k)}\partial_{\eta_{ab}^{(\ell)}}\mathcal{Z}_{\mathrm{out}}=\int\!\mathcal{D}\eta\;\partial_{\eta_{ab}^{(k)}}\bigl(\partial_{\eta_{ab}^{(\ell)}}\mathcal{Z}_{\mathrm{out}}\bigr).$ (124)
and exploiting the symmetry: $(\ell\leftrightarrow k)$, we finally arrive to
$\widehat{q}_{\ell k}=\frac{4\alpha}{q_{\ell k}}\sum_{a\leq b}\int\!\mathcal{D}\eta\,dy\;\partial_{\eta_{ab}^{(\ell)}}\mathcal{Z}_{\mathrm{out}}\;\partial_{\eta_{ab}^{(k)}}\ln\mathcal{Z}_{\mathrm{out}}.$ (125)
which, because $\partial_{\eta}\ln\mathcal{Z}_{\mathrm{out}}=(\partial_{\eta}\mathcal{Z}_{\mathrm{out}})/\mathcal{Z}_{\mathrm{out}}$, finally becomes

### v3

@@ C.3 Output Channel Computation
$I_{\rm out}(q)=\left[\int dy\int\prod_{u,\ell,a\leq b}dh^{(\ell)\;u}_{ab}\,P\left(\{h^{(\ell)\;u}_{ab}\}_{ab}\right)\prod_{u,\ell}P_{\rm out}\left(y\,\Bigm|\,\Bigl\{\frac{h^{(\ell)\;u}_{ab}}{\sqrt{2-\delta_{ab}}}\Bigr\}_{ab}\right)\right]^{n},$ (97)
@@
$\mathbb{E}[h^{(\ell)}_{ab}]=0,\qquad\operatorname{Cov}_{x^{\mu}}\left(h^{(\ell)\;u}_{a\leq b},h^{(k)\;v}_{c\leq d}\right)=\frac{1}{d}\left[2\;\delta_{ac}\delta_{bd}\right]\operatorname{Tr}\left(S^{u}_{\ell}S^{v}_{k}\right)=\left[2\;\delta_{ac}\delta_{bd}\right]Q^{uv}_{\ell k}.$ (98)
@@
$\frac{1}{2}\sum_{u,v=0}^{m}\sum_{a\leq b,c\leq d}\sum_{\ell,k}\biggl(h^{(\ell)\;u}_{ab}\biggr)\left(\Sigma_{h}^{-1}\right)^{uv,\ell k}_{ab,cd}\biggl(h^{(k)\;v}_{cd}\biggr).$ (99)
@@
$I_{\rm out}(q)=\left[\int dy\int\prod_{u,\ell,a\leq b}dh^{(\ell)\;u}_{ab}\,\exp\left(-\frac{1}{2}\sum_{u,v=0}^{m}\sum_{a,b,c,d}\sum_{\ell,k}h^{(\ell)\;u}_{ab}\left(\Sigma_{h}^{-1}\right)^{uv,\ell k}_{a\leq b,c\leq d}h^{(k)\;v}_{cd}\right)\prod_{u,\ell}P_{\rm out}\left(y\,\Bigm|\,\Bigl\{\frac{h^{(\ell)\;u}_{ab}}{\sqrt{2-\delta_{ab}}}\Bigr\}_{ab}\right)\right]^{n}.$
@@
$\left(\Sigma_{h}\right)^{uv,\ell k}_{a\leq b,c\leq d}=\left[2\;\delta_{ac}\delta_{bd}\right]\,Q^{uv}_{\ell k}=\left[2\;\delta_{ac}\delta_{bd}\right]\Bigl[(Q_{\ell k}-q_{\ell k})\,\delta_{uv}+q_{\ell k}\Bigr].$ (101)
Because of the Kronecker structure $\delta_{ac}\delta_{bd}$, the Gaussian law over all $\{h^{(\ell)u}_{ab}\}_{a\leq b,\ell,u}$ factorizes over token pairs $(a,b)$. Hence it is sufficient to treat one fixed pair $(a,b)$ and then take the product over $a\leq b$. For notational clarity in the next steps, we temporarily fix a pair $(a,b)$ and write $h^{(\ell)u}\equiv h^{(\ell)u}_{ab}$. For this pair, the covariance across replicas and layers reads
@@
$\mathbb{E}\!\left[h^{(\ell)u}h^{(k)v}\right]=2\Bigl[(Q_{\ell k}-q_{\ell k})\,\delta_{uv}+q_{\ell k}\Bigr],\qquad u,v=0,\dots,m,\ \ \ell,k=1,\dots,L.$ (102)
We now construct explicitly a family of Gaussian random variables with covariance (102). Introduce a shared Gaussian vector $\omega=(\omega^{(1)},\dots,\omega^{(L)})$ with mean zero and covariance
$\mathbb{E}\!\left[\omega^{(\ell)}\omega^{(k)}\right]=2q_{\ell k}.$ (103)
For each replica $u=0,\dots,m$, an independent Gaussian vector $\xi^{u}=(\xi^{(1)u},\dots,\xi^{(L)u})$ with mean zero and covariance
$\mathbb{E}\!\left[\xi^{(\ell)u}\xi^{(k)v}\right]=2(Q_{\ell k}-q_{\ell k})\,\delta_{uv}.$ (104)
$\omega$ is independent of all $\{\xi^{u}\}_{u=0}^{m}$. Define for each replica $u$ and layer $\ell$:
$h^{(\ell)u}:=\omega^{(\ell)}+\xi^{(\ell)u}.$ (105)
Then, for all $\ell,k$ and $u,v$,
$\mathbb{E}\!\left[h^{(\ell)u}h^{(k)v}\right]$ $=\mathbb{E}\!\left[(\omega^{(\ell)}+\xi^{(\ell)u})(\omega^{(k)}+\xi^{(k)v})\right]$
$=\mathbb{E}\!\left[\omega^{(\ell)}\omega^{(k)}\right]+\mathbb{E}\!\left[\xi^{(\ell)u}\xi^{(k)v}\right]$
$=2q_{\ell k}+2(Q_{\ell k}-q_{\ell k})\delta_{uv}=2\Bigl[(Q_{\ell k}-q_{\ell k})\delta_{uv}+q_{\ell k}\Bigr],$ (106)
which is exactly (102). Therefore the law of $\{h^{(\ell)u}\}_{\ell,u}$ is exactly the RS Gaussian law.
From (105), conditional on $\omega$ the replicas are independent and
$\{h^{(\ell)u}\}_{\ell=1}^{L}\ \Big|\ \omega\ \sim\ \mathcal{N}\!\left(\{\omega^{(\ell)}\}_{\ell=1}^{L},\ V\right),\qquad V^{(\ell k)}:=2(Q_{\ell k}-q_{\ell k}).$ (107)
Equivalently, for each $u$,
$p\!\left(\{h^{(\ell)u}\}_{\ell}\mid\omega\right)=\frac{1}{\sqrt{\det(2\pi V)}}\,\exp\!\left(-\frac{1}{2}\sum_{\ell=1}^{L}\sum_{k=1}^{L}\bigl(h^{(\ell)u}-\omega^{(\ell)}\bigr)\,[V^{-1}]_{\ell k}\,\bigl(h^{(k)u}-\omega^{(k)}\bigr)\right).$ (108)
Moreover $\omega\sim\mathcal{N}(0,2q)$, i.e.
$p(\omega)=\frac{1}{\sqrt{\det(2\pi\,2q)}}\,\exp\!\left(-\frac{1}{2}\sum_{\ell=1}^{L}\sum_{k=1}^{L}\omega^{(\ell)}\,[(2q)^{-1}]_{\ell k}\,\omega^{(k)}\right).$ (109)
By marginalization over $\omega$, the joint density of $\{h^{u}\}_{u=0}^{m}$ is therefore:
$p\!\left(\{h^{(\ell)u}\}_{\ell,u}\right)=\int d^{L}\omega\;p(\omega)\;\prod_{u=0}^{m}p\!\left(\{h^{(\ell)u}\}_{\ell}\mid\omega\right).$ (110)
We now express $\omega\sim\mathcal{N}(0,2q)$ through standard normals. Introduce i.i.d. auxiliary variables $\eta^{(1)},\dots,\eta^{(L)}\sim\mathcal{N}(0,1)$ and define
$\omega^{(\ell)}=\sum_{r=1}^{L}(\sqrt{2q})_{\ell r}\,\eta^{(r)},$ (111)
where $(\sqrt{2q})$ is any matrix square root such that, for all $\ell,k$,
$\sum_{r=1}^{L}(\sqrt{2q})_{\ell r}(\sqrt{2q})_{kr}=(2q)_{\ell k}.$ (112)
Then $\omega$ has covariance (103) since
$\mathbb{E}[\omega^{(\ell)}\omega^{(k)}]$ $=\sum_{r,s}(\sqrt{2q})_{\ell r}(\sqrt{2q})_{ks}\,\mathbb{E}[\eta^{(r)}\eta^{(s)}]=\sum_{r}(\sqrt{2q})_{\ell r}(\sqrt{2q})_{kr}=(2q)_{\ell k}.$ (113)
Consequently, for any function $F(\omega)$,
$\int d^{L}\omega\;\mathcal{N}(\omega;0,2q)\,F(\omega)=\int\prod_{r=1}^{L}\frac{d\eta^{(r)}}{\sqrt{2\pi}}e^{-(\eta^{(r)})^{2}/2}\;F\!\left(\left\{\sum_{r}(\sqrt{2q})_{\ell r}\eta^{(r)}\right\}_{\ell=1}^{L}\right).$ (114)
Restoring the token indices, for each $(a,b)$ we introduce independent $\eta^{(r)}_{ab}\sim\mathcal{N}(0,1)$ and define
$\omega_{ab}^{(\ell)}=\sum_{k=1}^{L}(\sqrt{2q})_{\ell k}\,\eta_{ab}^{(k)},\qquad V^{(\ell k)}=2(Q_{\ell k}-q_{\ell k}).$ (115)
Using (107)–(114), the Gaussian exponent for each replica $u$ and each token pair $(a,b)$ is therefore exactly
$\frac{1}{2}\sum_{\ell=1}^{L}\sum_{k=1}^{L}\bigl(h^{(\ell)\;u}_{ab}-\omega^{(\ell)}_{ab}\bigr)\;[V^{-1}]_{\ell k}\;\bigl(h^{(k)\;u}_{ab}-\omega^{(k)}_{ab}\bigr)\;-\frac{1}{2}\sum_{r=1}^{L}(\eta^{(r)}_{ab})^{2},$ (116)
which is the desired form with $\omega$ and $V$ identified as in (115).
Define the auxiliary measure
$\mathcal{D}\eta=\prod_{a\leq b}\prod_{\ell=1}^{L}\frac{d\eta_{ab}^{(\ell)}}{\sqrt{2\pi}}\,\exp\!\Bigl[-\frac{(\eta_{ab}^{(\ell)})^{2}}{2}\Bigr].$ (117)
For fixed $\eta$, the replicas are independent and identically distributed through the conditional Gaussian $\mathcal{N}(h_{ab};\omega_{ab},V)$ on layers $\ell$. Hence, introducing the single-replica output partition function
@@
$\mathcal{Z}_{\rm out}(y,\omega,V)=\int\prod_{a\leq b}d^{L}h_{ab}\;\mathcal{N}\!\left(h_{ab};\omega_{ab},V\right)\;P_{\rm out}\!\left(y\,\Bigm|\,\Bigl\{\frac{h^{(\ell)}_{ab}}{\sqrt{2-\delta_{ab}}}\Bigr\}_{a\leq b,\ \ell=1}^{L}\right),$ (118)
we can write
$I_{\rm out}(q)=\left[\int dy\int\mathcal{D}\eta\;\Bigl(\mathcal{Z}_{\rm out}(y,\omega,V)\Bigr)^{m+1}\right]^{n},$ (119)
where $\omega$ and $V$ are the functions of $(q,Q)$ defined in (115). Expanding Eq. (119) for small number of replicas $m\to 0$ and using $A^{m+1}=A\,e^{m\ln A}=A\,(1+m\ln A+o(m))$, the contribution to the free entropy is governed by
@@
We aim to compute $\partial_{q}I_{\rm out}(q)$ to finally reach the state equation in Eq. (86), namely:
@@
$\widehat{q}_{\ell k}=4\alpha\int dy\int\mathcal{D}\eta\;\bigl[1+\ln\mathcal{Z}_{\rm out}(y,\omega,V)\bigr]\;\frac{\partial\mathcal{Z}_{\rm out}(y,\omega,V)}{\partial q_{\ell k}}.$ (121)
It is convenient to rewrite the $\eta$-integral as an integral over $\omega$ itself. For each token pair $(a,b)$ we have $\omega_{ab}\sim\mathcal{N}(0,2q)$ with independent draws across $a\leq b$. Thus we may equivalently write
$I_{\rm out}(q)=\int dy\int\Bigl[\prod_{a\leq b}d^{L}\omega_{ab}\,\mathcal{N}(\omega_{ab};0,2q)\Bigr]\;\mathcal{Z}_{\rm out}(y,\omega,V)\,\ln\mathcal{Z}_{\rm out}(y,\omega,V),$ (122)
where $\omega=\{\omega_{ab}^{(\ell)}\}_{a\leq b,\ell}$. Now define for each $(a,b)$ and layer $\ell$:
$(g_{\rm out}(y,\omega,V))_{ab}^{(\ell)}:=\frac{\partial}{\partial\omega_{ab}^{(\ell)}}\ln\mathcal{Z}_{\rm out}(y,\omega,V).$ (123)
In (122), $q$ appears in: (i) in the Gaussian measure $\omega_{ab}\sim\mathcal{N}(0,2q)$, and (ii) in V=2(Qq)V=2(Q-q) inside $\mathcal{Z}_{\rm out}$. Let
$F(\omega,q):=\int dy\;\mathcal{Z}_{\rm out}(y,\omega,V)\,\ln\mathcal{Z}_{\rm out}(y,\omega,V).$
Then
$I_{\rm out}(q)=\int\Bigl[\prod_{a\leq b}d^{L}\omega_{ab}\,\mathcal{N}(\omega_{ab};0,2q)\Bigr]\;F(\omega,q).$ (124)
Now fix a single pair $(a,b)$, under $\omega_{ab}\sim\mathcal{N}(0,\Sigma)$ with $\Sigma=2q$, we get the Gaussian identity, for any smooth $G(\omega_{ab})$:
$\frac{\partial}{\partial\Sigma_{\ell k}}\mathbb{E}_{\omega_{ab}\sim\mathcal{N}(0,\Sigma)}[G(\omega_{ab})]=\frac{1}{2}\,\mathbb{E}_{\omega_{ab}}\!\left[\frac{\partial^{2}G(\omega_{ab})}{\partial\omega^{(\ell)}_{ab}\,\partial\omega^{(k)}_{ab}}\right$]. (125)
Since $\Sigma=2q$, we have $\partial/\partial q_{\ell k}=2\,\partial/\partial\Sigma_{\ell k}$, hence
$\frac{\partial}{\partial q_{\ell k}}\mathbb{E}_{\omega_{ab}\sim\mathcal{N}(0,2q)}[G(\omega_{ab})]=\mathbb{E}_{\omega_{ab}}\!\left[\frac{\partial^{2}G(\omega_{ab})}{\partial\omega^{(\ell)}_{ab}\,\partial\omega^{(k)}_{ab}}\right$]. (126)
Applying (126) to (124) (and summing over independent pairs) yields